OpenStudy (anonymous):

The school is building a basketball court for its students. The length of the basketball court is 31 1/8 feet. The width of the basketball court is 25 1/2 feet. What is the perimeter of the basketball court?

OpenStudy (lgbasallote):

perimeter = length + length + width + width \[\implies 32 \frac 18 + 31\frac 18 + 25 \frac 12 + 25 \frac 12\] does that help?

OpenStudy (lgbasallote):

that should be 31 1/8 not 32

OpenStudy (anonymous):

ok

OpenStudy (lgbasallote):

k.

OpenStudy (anonymous):

thks

OpenStudy (jiteshmeghwal9):

@shivraj perimeter of a rectangle=2(l+b) in easy way but @lgbasallote is also correct:)

OpenStudy (lgbasallote):

heh i dont believe in "easy ways" i just believe in ways lol

OpenStudy (jiteshmeghwal9):

HaHaHa!

OpenStudy (anonymous):

ok thnks